OpenStudy (anonymous):

integrate 3u^2du/(1+u) from 0 to 1

Parth (parthkohli):

\[\int_{0}^{1} {3u^2du \over 1 + u}\]Is that the question?

OpenStudy (anonymous):

yes

Parth (parthkohli):

OK, can you first perform the indefinite integration?

OpenStudy (anonymous):

Im having trouble setting that up

OpenStudy (anonymous):

I dont see how to break it into partial fractions

Parth (parthkohli):

I have to go to sleep, sorry for the inconvenience... :(

OpenStudy (anonymous):

3u² = 3 ( u² - 1 + 1 ) = 3 ( u -1) ( u+1) + 3

OpenStudy (anonymous):

-> 3 ( u -1 ) + 3 / ( u +1)

myininaya (myininaya):

Yes! You can do long division since the degree on top is greater than or equal to the degree of the polynomial on the bottom. Or do what @Chlorophyll did which is a really really cute way.

OpenStudy (anonymous):

@myininaya Thanks for nice compliment. It's actually the lazy way :")